in the following a set V, a field F, which is either R or C, and operations of addition + and scalar multiplication . , are given for alpha element of F and x element of V, we write their multiplication alpha × x as alpha • x, cheak whether V is a vector space over F, with these operations .
1
Expert's answer
2020-10-28T19:33:21-0400

As I understand it, we want to check 10 axioms of a vector space.

1) AC Additive Closure


"\text{if}\quad \vec{u},\vec{v}\in V, \quad\text{then}\quad \left(\vec{u}+\vec{v}\right)\in V"

2) SC Scalar Closure


"\text{if}\quad\alpha\in\mathbb{C}\quad\text{and}\quad \vec{u}\in V,\quad\text{then}\quad(\alpha\cdot\vec{u})\in V"

3) C Commutativity


"\text{if}\quad \vec{u},\vec{v}\in V, \quad\text{then}\quad \vec{u}+\vec{v}=\vec{v}+\vec{u}"

4) AA Additive Associativity


"\text{if}\quad \vec{u},\vec{v},\vec{w}\in V, \quad\text{then}\quad \left(\vec{u}+\vec{v}\right)+\vec{w}=\vec{v}+\left(\vec{u}+\vec{w}\right)"

5) Z Zero Vector


"\text{There is a vector},\,\,\, \vec{0},\,\,\,\text{called the zero vector, such that}\\[0.3cm]\,\,\,\vec{u}+\vec{0}=\vec{u}\,\,\,\text{for all}\,\,\,\vec{u}\in V"

6) AI Additive Inverses


"\text{If}\,\,\,\vec{u}\in V,\,\,\,\text{then there exists a vector}\,\,\,(−\vec{u})\in V\,\,\,\text{so that}\\[0.3cm]
\vec{u}+(−\vec{u})=\vec{0}."

7) SMA Scalar Multiplication Associativity


"\text{If} \,\,\,\alpha, \beta\in\mathbb{C}\,\,\,\text{and}\,\,\,\vec{u}\in V, \,\,\,\text{then}\\[0.3cm] 
\alpha\cdot(\beta\cdot\vec{u})=(\alpha\beta)\cdot\vec{u}"

8) DVA Distributivity across Vector Addition


"\text{If}\,\,\,\alpha\in\mathbb{C}\,\,\,\text{and}\,\,\,\vec{u},\vec{v}\in V,\,\,\,\text{then}\\[0.3cm] 
\alpha\cdot(\vec{u}+\vec{v})=\alpha\cdot\vec{u}+\alpha\cdot\vec{v}"

9) DSA Distributivity across Scalar Addition


"\text{If}\,\,\,\alpha,\beta\in\mathbb{C}\,\,\,\text{and}\,\,\,\vec{u}\in V,\text{then}\\[0.3cm] 
(\alpha+\beta)\cdot\vec{u}=\alpha\cdot\vec{u}+\beta\cdot\vec{u}"

10) O One


"\text{If}\,\,\,\vec{u}\in V,\,\,\,\text{then}\,\,\, 1\cdot\vec{u}=\vec{u}."